# Mapple: A Domain-Specific Language for Mapping Distributed Heterogeneous Parallel Programs

### v1

@@ 4.3. Algorithm Implementation and Complexity Analysis
We now analyze the complexity of the search-based algorithm. Since the algorithm exhaustively enumerates all valid factorization strategies, its complexity is determined by the size of the search space. As shown in the earlier example, enumerating non-negative integer solutions to the equation $x_{1}+x_{2}+x_{3}=4$ is equivalent to counting integer partitions with repetition, which can be mapped to the problem of selecting two dividers among six gaps. This corresponds to the number of combinations $\binom{6}{2}=15$.
In general, for a processor count with prime factorization $d_{i}=p_{1}^{a_{1}}\cdot p_{2}^{a_{2}}\cdots p_{t}^{a_{t}}$, the total number of factorization strategies is:
$\prod_{j=1}^{t}\binom{a_{j}+k-1}{k-1}$
Each term in the product counts the number of ways to distribute the $a_{j}$ copies of prime $p_{j}$ across $k$ dimensions. In practice, the exponents $a_{j}$ are usually small (typically less than 10), and the number of dimensions $k$ is often no greater than 3. As a result, the search space remains small, and the algorithm is computationally efficient in real-world scenarios.
We believe that enumeration is necessary to guarantee an optimal solution. Consider a greedy algorithm that, at each step, assigns prime factors of $d_{i}=72$ (i.e., $2,2,2,3,3$) to dimensions in a way that minimizes the maximum difference between elements of the workload vector. Given an iteration space of $(l_{1},l_{2})=(8,9)$, this greedy strategy produces a suboptimal workload vector of $\left(\frac{4}{3},\frac{3}{4}\right)$, which results in imbalanced partitioning. In contrast, our search-based algorithm identifies the optimal factorization that yields a perfectly balanced workload vector of $(1,1)$.
@@ 5. Implementation
We describe the challenges and implementation of translating Mapple into task-based runtime systems. These systems employ a highly asynchronous and pipelined execution model, where tasks progress through multiple stages to maximize throughput. Mapping decisions—determining where tasks and data are placed—must be made at various points along this pipeline and are exposed via many distinct callback functions, each tied to a specific stage in a task’s lifecycle. This results in a fragmented and low-level interface that mirrors the internal structure of the runtime rather than providing an intuitive programming model for developers.
@@ 5.1. Execution Semantics of the Runtime System
Task-based runtime systems are designed to maximize throughput through an asynchronous, pipelined execution model, where tasks advance through several distinct stages during their lifetime. To illustrate the low-level nature of the mapping interface, we present the execution semantics of such systems. This exposes a key source of complexity: the interface is tightly coupled to internal execution stages and operates at a low level of abstraction. Consequently, a central challenge is to design a high-level, centralized mapping DSL that abstracts away low-level details while still capturing critical mapping decisions.
Before presenting the semantics, we first describe the key relationships between tasks that govern how and when they execute. Consider the following simple program:
task f(a, b) {
g(a); // writes a
h(b); // writes b
k(a, b); // reads a, b and writes b
@@
[…]
Figure 10 defines the structures used in the execution semantics. The execution state consists of two vectors of $n$ queues—one for tasks enqueued on each node, and one for tasks that have been mapped. These represent two of the four stages in a task’s lifetime; the others are when a task is launched and when it is executed. A task progresses through these stages as follows:
  • Enqueued: A task is enqueued once its parent and all sibling predecessors have been enqueued, preserving program order and control dependencies.
  • Mapped: A task is mapped after all sibling predecessors it depends on have been mapped, ensuring their processor locations are known for scheduling data movement.
  • Launched: A task can launch after all dependence predecessors have executed, ensuring their effects are complete and visible.
  • Executed: A task is considered executed once it finishes running and all of its children have also completed, since their effects contribute to the parent’s semantics.
$\begin{array}[]{rl}\mbox{\em State}&\\ \mbox{Execution State }S&::=N[n]\ast N[n]\\ \mbox{Execution Log }L&::=e~{}\mbox{\bf set}\\ \mbox{Execution Log Entry }e&::=\mbox{enqueued}(t)\mid\mbox{mapped}(t,p)\mid\mbox{launched}(t,p)\mid\mbox{executed}(t,p)\\ \mbox{Node Queues }N&::=(t~{}\mbox{\bf queue})\\ \mbox{Processor }p&::=id\\[14.45377pt] \mbox{\em Tasks}&\\ \mbox{Point }P&::=(i_{1},\ldots,i_{n})\\ \mbox{Task }t&::=\mbox{Index}(id,P~{}\mbox{\bf set})\\[14.45377pt] \mbox{\em Relations}&\\ \mbox{Task Tree Relation }r&::=\mbox{parent}(t_{p},t_{c})\mid\mbox{task}(t)\\ \mbox{Task Tree }T&::=r~{}\mbox{\bf set}\\ \mbox{Dependence Relation }d&::=t\leq t^{\prime}\\ \mbox{Dependencies }D&::=d~{}\mbox{\bf set}\\ \mbox{Sibling Relation }w&::=t\preceq t^{\prime}\\ \mbox{Siblings }W&::=w~{}\mbox{\bf set}\\ \end{array}$
@@
$\frac{\begin{array}[]{l}task(t)\in T\\ t=\text{Index}(id,P)\\ \text{enqueued}(t)\notin L\\ \text{parent}(t_{p},t)\in T\\ \text{launched}(t_{p},p)\in L\\ \forall t_{s}:\text{parent}(t_{p},t_{s})\in T\wedge t_{s}\preceq t,\,\text{enqueued}(t_{s})\in L\end{array}}{L;(E,M)\xrightarrow{T;D;W}L\cup\{\text{enqueued}(t)\};(E+_{p}t,M)}$ [Enqueue]
$\frac{\begin{array}[]{l}t=\text{Index}(id,P)\\ \text{SHARD}(t)=\text{distribute}((t_{1},p_{1}),(t_{2},p_{2}))\\ t=(id,P)\\ t_{1}=(id_{1},P_{1})\\ t_{2}=(id_{2},P_{2})\\ id_{1},id_{2}\text{ fresh (not in $T,E,M,L$)}\\ P=P_{1}\cup P_{2}\\ E^{\prime}=E-_{i}t\end{array}}{L;(E,M)\xrightarrow{T;D;W}L;((E^{\prime}+_{p_{1}}t_{1})+_{p_{2}}t_{2},M)}$ [Distribute]
$\frac{\begin{array}[]{l}\text{SHARD}(t)=\text{local}(t)\\ E^{\prime}=E-_{p}t\end{array}}{L;(E,M)\xrightarrow{T;D;W}L;(E^{\prime},M+_{p}t)}$ [Local]
$\frac{\begin{array}[]{l}\forall t^{\prime}:t^{\prime}\leq t,\,\text{mapped}(t^{\prime},\_)\in L\\ \text{MAP}(t)=p\\ M^{\prime}=M-_{i}t\end{array}}{L;(E,M)\xrightarrow{T;D;W}L\cup\{\text{mapped}(t,p)\};(E,M^{\prime})}$ [Map]
$\frac{\begin{array}[]{l}\forall t^{\prime}:t^{\prime}\leq t,\,\text{executed}(t^{\prime},\_)\in L\\ \text{mapped}(t,p)\in L\end{array}}{L;(E,M)\xrightarrow{T;D;W}L\cup\{\text{launched}(t,p)\};(E,M)}$ [Launch]
$\frac{\begin{array}[]{l}\forall t^{\prime}:t^{\prime}\leq t,\,\text{executed}(t^{\prime},\_)\in L\\ \forall c:\text{parent}(t,c),\,\text{executed}(c,\_)\in L\\ \text{mapped}(t,p)\in L\end{array}}{L;(E,M)\xrightarrow{T;D;W}L\cup\{\text{executed}(t,p)\};(E,M)}$ [Execute]
Figure 11. Transition relation judgments.
@@
$L;(E,M)\xrightarrow{T;D;W}L^{\prime};(E^{\prime},M^{\prime})$
which should be read as: given a task tree $T$, a dependence relation $D$, and a sibling predecessor relation $W$, the current execution log $L$ and execution state $(E,M)$, representing the sets of enqueued and mapped tasks on each node, transition to a new execution log $L^{\prime}$ and new state $(E^{\prime},M^{\prime})$. The rules rely on two user-supplied mapping functions, SHARD and MAP, with the following signatures:
SHARD: $\begin{array}[]{rl}\mbox{SHARD: }&t\rightarrow\mbox{distribute}((t\ast p)\ast(t\ast p))+\mbox{local}(t)\\ \mbox{MAP: }&t\rightarrow p\end{array}$
The SHARD function takes an index task and either partitions it into smaller tasks for distribution across nodes or assigns it to the local node. MAP then selects a specific processor within the target node for execution. Although these are two separate callback functions in the low-level interface, Mapple unifies them by interpreting both as part of a single index transformation from iteration space to processor space.
Figure 11 formalizes the rules described above as execution judgments. The rules use a new operator $+_{i}$, which adds a task to the $i$th queue of a vector of queues:
@@
We briefly summarize each rule. [Enqueue] places a task on the same node as its parent once the parent has launched and all sibling predecessors are enqueued. [Distribute] partitions an index task into independent sub-tasks for mapping. [Local] selects the target node based on the user-defined SHARD function. [Map] assigns the task to a specific processor within that node using the MAP function. [Launch] and [Execute] mark the start and completion of execution and involve no further mapping decisions.
The semantics of task execution underscore the low-level nature of mapping in task-based systems. To achieve high performance, tasks progress through multiple pipelined stages, each requiring user-defined mapping decisions. For brevity, we focus on Shard and MAP, though in practice, 19 distinct callback APIs are invoked across a task’s lifetime. This design fragments mapping logic across numerous low-level callbacks, making the APIs complex and opaque to most application developers.
@@ 5.2. Translation
The translation poses two key challenges. First, as discussed in Section 5.1, mapping logic is spread across 19 callback functions triggered at different stages of a task’s lifetime. While this mirrors the runtime’s internal pipeline, it forces developers to work with a scattered and unintuitive interface (Figure 1). Second, the low-level API lacks abstraction over the machine architecture—there is no logical view of the hardware, nor support for aligning mapping decisions with loop structures or restructuring index spaces to facilitate coordination.
We implement Mapple by targeting Legion’s programmatic mapping interface, comprising roughly 7,000 lines of C++ code—much of it dedicated to handling low-level runtime details. Mapple overrides a set of callback functions invoked at different stages of the task execution pipeline to decide task placement. The two core overrides, corresponding to the user-defined SHARD and MAP functions, are shown in Figure 1(b) and explained in Section 5.1.
In Mapple, the machine’s logical view is defined with m = Machine(GPU), and users apply index transformations using provided primitives. The runtime invokes the SHARD function to assign a task to a node, followed by MAP to select a processor on that node. Both are implemented by interpreting user-defined mapping functions written in Mapple, evaluated per iteration point. The core translation computes the node (for SHARD) and processor (for MAP) identifiers via a sequence of invertible transformations over the processor space. As defined in Figure 6, these transformations yield a pair of unsigned integers representing coordinates in the original 2D processor space specified by Machine. These coordinates are then returned as the node and processor identifiers.
@@ 6. Results
Benchmark.
Our approach is evaluated using a diverse set of nine benchmarks selected to demonstrate the generality and effectiveness of Mapple across representative distributed workloads. Matrix multiplication is one of the most fundamental and heavily studied problems in parallel computing, serving as a cornerstone for many scientific applications. It is also a classic stress test for distributed systems, due to the need to balance computation and communication across complex processor hierarchies. The first six benchmarks therefore, focus on advanced parallel matrix multiplication algorithms: Cannon’s (Cannon, 1969), SUMMA (Van De Geijn and Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik and Demmel, 2011), and COSMA (Kwasniewski et al., 2019). They are divided into 2D and non-2D categories. The 2D algorithms, such as Cannon’s, PUMMA, and SUMMA, enhance communication efficiency by partitioning matrices into 2D tiles and mapping them onto processor space, utilizing techniques like pipelining. Conversely, non-2D algorithms like Johnson’s, Solomonik’s, and COSMA, employ 3D partitioning or optimize processor grids to minimize communication overhead and boost performance.
The remaining three benchmarks are drawn from scientific simulation workloads, which represent a different class of distributed applications with distinct computation models and data access patterns. Circuit (Bauer et al., 2012) simulates electrical circuit behavior by modeling currents and voltages across interconnected nodes and wires. Stencil (Van der Wijngaart and Mattson, 2014) operates on a 2D grid, updating each point’s value based on a stencil pattern derived from its neighbors. Pennant (Ferenbaugh, 2015) models unstructured mesh Lagrangian staggered-grid hydrodynamics, essential for simulating compressible flow.
Together, these benchmarks span a wide spectrum of distributed computation patterns. Demonstrating strong results on these diverse workloads provides convincing evidence of Mapple’s ability to express complex mapping strategies and deliver high performance across application domains.
@@ 6.1. Lines of Code Reduction
We show the lines of code comparison between Mapple and C++ mappers in Table 1. For each mapper, we count the number of non-blank, non-comment lines of code written for the Mapple mapper and the C++ mapper. Using Mapple to implement mappers leads to a reduction in lines of code from 406 to 29 on average, i.e., a 14$\times$ reduction across all benchmarks. In Figure 12, we present example functions utilized by the mappers for distributed matrix-multiplication algorithms. The DSL effectively captures all the mapping decisions made by the C++ mappers in a concise manner.
We manually verify that both approaches make identical mapping decisions and achieve matching performance (i.e., identical throughput), indicating that any overhead introduced by Mapple is negligible. This confirms the fidelity of our Mapple mappers when compared against prior low-level C++ implementations (Yadav et al., 2022a; Bauer et al., 2012).
Table 1. Lines of Code (LoC) comparison between DSL and C++ mappers. The DSL achieves a 14$\times$ average reduction in LoC while maintaining the same performance as the C++ mappers.
@@
LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14$\times$
Figure 12. Example functions used by the mappers for the distributed matrix-multiplication algorithms.

### v2

@@ 4.3. Algorithm Implementation and Complexity Analysis
We now analyze the complexity of the search-based algorithm. Since the algorithm exhaustively enumerates all valid factorization strategies, its complexity is determined by the size of the search space. As shown in the earlier example, enumerating non-negative integer solutions to the equation $x_{1}+x_{2}+x_{3}=4$ is equivalent to counting integer partitions with repetition, which can be mapped to the problem of selecting two dividers among six gaps. This corresponds to the number of combinations $\binom{6}{2}=15$. In general, for a processor count with prime factorization $d_{i}=p_{1}^{a_{1}}\cdot p_{2}^{a_{2}}\cdots p_{t}^{a_{t}}$, the total number of factorization strategies is $\prod_{j=1}^{t}\binom{a_{j}+k-1}{k-1}$. Each term in the product counts the number of ways to distribute the $a_{j}$ copies of prime $p_{j}$ across $k$ dimensions. In practice, the exponents $a_{j}$ are usually small (typically less than 10), and the number of dimensions $k$ is often no greater than 3. As a result, the search space remains small, and the algorithm is computationally efficient in real-world scenarios.
We believe that enumeration is necessary to guarantee an optimal solution. Consider a greedy algorithm that, at each step, assigns prime factors of $d_{i}=72$ (i.e., $2,2,2,3,3$) to dimensions in a way that minimizes the maximum difference between elements of the workload vector. Given a task space of $(l_{1},l_{2})=(8,9)$, this greedy strategy produces a suboptimal workload vector of $\left(\frac{4}{3},\frac{3}{4}\right)$, which results in imbalanced partitioning. In contrast, our search-based algorithm identifies the optimal factorization that yields a perfectly balanced workload vector of $(1,1)$.
@@ 5. Implementation
We describe the translation of Mapple to the mapping interface of task-based runtime systems. These systems employ a highly asynchronous and pipelined execution model, where tasks progress through multiple stages to maximize throughput. Mapping decisions, which determine where tasks and data are placed, must be made at various points along this pipeline and are exposed via many distinct callback functions, each tied to a specific stage in a task’s lifecycle. This results in a fragmented and low-level interface that mirrors the internal structure of the runtime rather than providing an intuitive programming model for developers.
@@ 5.1. Execution Semantics of the Runtime System
Task-based runtime systems maximize throughput through an asynchronous, pipelined execution model, where a task advances through several distinct stages during its lifetime. To illustrate the low-level nature of the mapping interface, we present a much simplified version of the execution semantics of the Legion runtime system (bauer2012legion), focusing on the portion of the pipeline that determines the processor on which a task runs. Our semantics makes explicit a key source of complexity: the interface is tightly coupled to internal execution stages, where what is conceptually a single decision—choosing a processor for a task—is split across multiple pipeline stages at a low level of abstraction. The same low-level complexity exists in frameworks like StarPU (augonnet2010starpu), Charm++ (kale1993charm++), and HPX (kaiser2014hpx), which expose mapping through asynchronous callbacks and fragmented APIs. In contrast, Mapple provides a high-level DSL that centralizes this logic while still providing the same flexibility.
Figure 11 defines the structures used in the execution semantics. A task space is a set of multidimensional points, each representing a single point task to be executed. Working with sets of tasks is important for the efficiency of the runtime system itself, especially in the early stages of the task pipeline, where runtime operations can be done in constant time on a set rather than in linear time on each individual point task in the set.
The execution context consists of a set of tasks TT [^1] One major simplification is that we ignore dependencies between tasks; dependencies are not needed to illustrate the pipeline for choosing a task’s processor., a log of actions on tasks, and two vectors of $n$ queues, one queue per node of the machine. [^2] We also do not model nodes, using only processors, but in a full description a set of processors is associated with each node. When a task space is first enqueued in the runtime system, it goes into a distribution queue, from where its set of points may be subdivided and sent to other nodes for further processing. Eventually (perhaps after multiple recursive subdivisions of the set of tasks), a task space is moved to a mapping queue on a particular processor, where the assignment of the task space to a (potentially different) processor for execution will be made. This structure, where task spaces are subdivided and spread out to multiple processors for mapping, parallelizes the mapping process, preventing a single processor from becoming a runtime bottleneck. Once a task space is mapped, its point tasks are enqueued for execution on its assigned processor; because our focus is on the mapping process, we do not illustrate task execution here.
@@
$\frac{\begin{array}[]{l}t\in T\\ \text{enqueued}(t)\notin L\\ \end{array}}{L,E,M\xrightarrow{T}L\cup\{\text{enqueued}(t)\},E+_{1}t,M}$ [Enqueue]
$\frac{\begin{array}[]{l}\text{SHARD}(t)=\text{distribute}((t_{1},p_{1}),(t_{2},p_{2}))\\ t=(id,P),\ t_{1}=(id_{1},P_{1}),\ t_{2}=(id_{2},P_{2})\\ id_{1},id_{2}\text{ fresh (not in $T,E,M,L$)}\\ P=P_{1}\cup P_{2}\\ \end{array}}{L,E,M\xrightarrow{T}L,((E-_{p}t)+_{p_{1}}t_{1})+_{p_{2}}t_{2},M}$ [Distribute]
$\frac{\begin{array}[]{l}\text{SHARD}(t)=\text{local}(t)\\ t=(id,P)\\ E^{\prime}=E-_{i}t\end{array}}{L,E,M\xrightarrow{T}L,E^{\prime},M+_{i}t}$ [Local]
$\frac{\begin{array}[]{l}\text{SLICE}(t)=p\\ M^{\prime}=M-_{i}t\end{array}}{L,E,M\xrightarrow{T}L\cup\{\text{mapped}(t,p)\},E,M^{\prime}}$ [Map]
Figure 11. Formalization of the mapping pipeline.
@@
$L,E,M\xrightarrow{T}L^{\prime},E^{\prime},M^{\prime}$
which should be read as: given a set of task spaces $T$, the current execution log $L$ and vectors of queues $E,M$, transition to a new execution log $L^{\prime}$ and new vectors of queues $E^{\prime},M^{\prime}$. The rules rely on two user-supplied mapping functions, SHARD and SLICE, with the following signatures:
$\begin{array}[]{rl}\mbox{SHARD: }&t\rightarrow\mbox{distribute}((t\ast p)\ast(t\ast p))+\mbox{local}(t)\\ \mbox{SLICE: }&t\rightarrow p\end{array}$
The SHARD function takes a task space and either partitions it into smaller task spaces for distribution across nodes or assigns it to the local node. SLICE then selects a specific processor within the target node for execution. Although these are two separate callback functions in the low-level interface, Mapple unifies them by interpreting both as part of a single index transformation from task space to processor space.
Figure 11 formalizes the rules described above as execution judgments. The rules use an operator $_{i}$, which adds a task space to the $i$th queue of a vector of queues:
@@
Note that dequeue is only defined if the task space $t$ is at the head of the selected queue.
We briefly summarize each rule. [Enqueue] submits a task space to the runtime system on node 1.333A simplification; in actual implementations, tasks can begin on any node.[Distribute] partitions a task space into smaller task spaces for mapping. [Local] selects the node where a task space’s mapping decision will be made using SHARD. [Map] assigns a task space to a specific processor within that node using the SLICE function.
The semantics of task execution underscore the low-level nature of mapping in task-based systems. To achieve high performance, task spaces progress through multiple pipelined stages, each requiring user-defined mapping decisions. For brevity, we focus on SHARD and SLICE, though in practice, 19 distinct callback APIs are invoked across a task’s lifetime. This design fragments mapping logic across numerous low-level callbacks, making the APIs complex and opaque to most application developers.
@@ 5.2. Translation
The translation of Mapple code to the low-level runtime interface poses two key challenges. First, as discussed in Section 5.1, Legion’s mapping logic is spread across many callback functions triggered at different stages of a task’s lifetime. While this structure mirrors the runtime’s internal pipeline, it forces developers to work with a scattered and unintuitive interface (Figure 1). Second, the low-level API lacks abstraction over the machine architecture, i.e., there is no logical view of the hardware, nor support for aligning mapping decisions with task spaces.
We implement Mapple by targeting Legion’s programmatic mapping interface, comprising roughly 7,000 lines of C++ code—much of it dedicated to handling low-level runtime details. Mapple overrides a set of callback functions invoked at different stages of the task execution pipeline to decide task placement. The two core overrides, corresponding to the user-defined SHARD and SLICE functions, are illustrated in Figure 1 and explained in Section 5.1.
We now give a concrete example using the block1D_x distribution (Figure 7). Let the processor space m = Machine(GPU) have shape $(n,g)$, representing $n$ nodes and $g$ GPUs per node. The mapper applies $m_{1}=m.\mathtt{merge}(0,1).\mathtt{split}(0,ng)$, which reshapes the processor space to $(ng,1)$. For an index point $(i_{x},i_{y})$ in a task space of size $(L_{x},L_{y})$, the block1D_x function returns point $(\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor,0)$ in $m_{1}$. The Mapple implementation then converts this point in $m_{1}$ back to a point in the original 2D processor space $m$ by applying the processor space transformations in reverse order (i.e., first inverting $\mathtt{split}(0,ng)$ and then inverting $\mathtt{merge}(0,1)$). Thus, the $m_{1}$ index $(\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor,0)$ maps back to a point in $m$ as $(\text{node\_id},\text{gpu\_id})=(a\bmod n,\lfloor a/n\rfloor)$ where $a=\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor$. In the original space of $m$, node_id identifies the node and gpu_id specifies the GPU within that node. The node index is used by SHARD, applying either [Distribute] or [Local] depending on whether the target node is local, while the GPU index is used by SLICE under the [Map] rule.
@@ 6. Results
Our approach is evaluated on a suite of nine benchmarks chosen to demonstrate the generality and effectiveness of Mapple across representative distributed workloads. We use the published artifact repositories (yadav2022distal; bauer2012legion) and their corresponding problem sizes, where the mappers were developed by experts and tuned for the same machine, representing state-of-the-art implementations.
The first three benchmarks are drawn from scientific simulation workloads. Circuit (bauer2012legion) simulates electrical circuit behavior by modeling currents and voltages across interconnected nodes and wires. Stencil (stencilvan2014parallel) operates on a 2D grid, updating each point’s value based on a stencil pattern derived from its neighbors. Pennant (ferenbaugh2015pennant) models unstructured mesh Lagrangian staggered-grid hydrodynamics, essential for simulating compressible flow.
The remaining six benchmarks are parallel matrix multiplication algorithms, an extensively studied problem in parallel computing that underpins many scientific applications. Matrix multiplication also serves as a classic stress test for distributed systems, as it requires carefully balancing computation and communication across multiple processors. They matrix multiplication algorithms we consider are Cannon’s (cannon1969cellular), SUMMA (van1997summa), PUMMA (choi1994pumma), Johnson’s (johnsonagarwal1995three), Solomonik’s (solomonik2011communication), and COSMA (cosmakwasniewski2019red). They are divided into 2D and non-2D categories. The 2D algorithms, such as Cannon’s, PUMMA, and SUMMA, enhance communication efficiency by partitioning matrices into 2D tiles and mapping them onto the processor space, utilizing techniques like pipelining. Conversely, non-2D algorithms like Johnson’s, Solomonik’s, and COSMA, employ 3D partitioning or optimize processor grids to minimize communication overhead and boost performance.
@@ 6.1. Lines of Code Reduction
[…]
We manually verify that both approaches produce identical mapping decisions and matching performance (i.e., identical throughput), showing that any overhead introduced by Mapple is negligible. This confirms the correctness and efficiency of our Mapple mappers relative to the original low-level C++ implementations in the published artifact repositories (yadav2022distal; bauer2012legion).
We note that the application code is substantially larger than the mapping code and remains unchanged. The three scientific computing applications, Circuit, Stencil, and Pennant, are implemented in Regent (slaughter2015regent) with 776, 825, and 3.6K lines, respectively, and the six distributed matrix multiplication benchmarks each contain about 7.5K lines of Legion C++ code.
[…]
@@
[…]